THE RESULTANT OF TWO FORCES WHICH ARE OF MAGNITUDE 130N AND 90N IS 150N FIND THE ANGLE BETWEEN THE TWO FORCES AND FIND THE ANGLE BETWEEN THE 150N AND 130N FORCE
1
Expert's answer
2019-10-23T09:42:29-0400


The law of cosines gives


"150^2=90^2+130^2-2\times 90\times 130\cos(180^{\circ}-\alpha)"

So


"\cos(180^{\circ}-\alpha)=0.1068"

Therefore

"\alpha=96^{\circ}"

The sine rule gives

"\frac{150}{\sin(180^\circ-\alpha)}=\frac{90}{\sin(\theta)}"

So

"\sin(\theta)=\frac{90}{150}\times\sin(84^\circ)=0.5967"

Finally


"\theta=37^{\circ}"